How do you find the radius and area of the circle. The circumference of a circle is 4 inches.
Answer by lenny460(1073)

How do you find the radius and area of the circle. The circumference of a circle is 4 inches.

Circumference of a circle = pi * Diameter
pi = 3.14
Diameter = 2 * radius
Therefore:
radius = circumference / (2 * pi)
4 / (2 * 3.14) = 4 / 6.28 = 0.63
radius = 0.63 inches

Area of a circle = pi * r^2
3.14 * 0.63^2
0.63 * 0.63 = 0.4044
3.14 * 0.4044 = 1.27011
Area of the circle = 1.27 inches squared.